078898:

Molly's alarm clock fell to the floor and the glass face was broken. Molly measured the glass face so that she could order a new one. The radius of the circle-shaped glass face is 6 centimeters. What is the approximate area of the glass face? a. 113.04 cm ,B.37.68 cm , C. 18.84 cm, D.452.16 cm. which one is it?

Shadow:

\[A_{circle} = \pi r^2\] \[A_{circle} = \pi \times (6)^2 = \pi \times 36 = 113.1 cm^2\]

078898:

thanks